OpenStudy (anonymous):

A sporting good store receives a shipment of 124 golf bags. The shipment includes 2 types of bags, full-size and collapsible. The full-size bags cost $38.50 each. The collapsible bags cost $22.50 each. The bill for the shipment is $3430. How many of each type of golf bag are in the shipment? I need to define variables and write equations. I already have variables so how do I find the constraints?

jimthompson5910 (jim_thompson5910):

Let f = # of full-size bags c = # of collapsible bags "A sporting good store receives a shipment of 124 golf bags" which means f+c = 124 "The full-size bags cost $38.50 each. The collapsible bags cost $22.50 each. The bill for the shipment is $3430" telling us 38.50f + 22.50c = 3430

jimthompson5910 (jim_thompson5910):

So you have these two equations f+c = 124 38.50f + 22.50c = 3430 do you know how to solve from here?

OpenStudy (anonymous):

not really

jimthompson5910 (jim_thompson5910):

Solve f+c = 124 for one of the variables I'll pick c f+c = 124 c = 124-f

jimthompson5910 (jim_thompson5910):

Then you plug c = 124-f into the second equation 38.50f + 22.50c = 3430 38.50f + 22.50(124-f) = 3430 now solve for f

OpenStudy (anonymous):

ok....thanks a lot!

jimthompson5910 (jim_thompson5910):

once you find f, you can find c by plugging it into c = 124-f

OpenStudy (anonymous):

I'm lost on how to solve for f...

jimthompson5910 (jim_thompson5910):

I'll get you started 38.50f + 22.50(124-f) = 3430 38.50f + 22.50(124)+22.50(-f) = 3430 38.50f + 2790 - 22.50f = 3430 16f + 2790 = 3430 f = ??

OpenStudy (anonymous):

is it like this: 38.50f = 22.50(124-f)=3430?

OpenStudy (anonymous):

Oh, I see you wrote that

OpenStudy (anonymous):

I got my x and y number, how would I graph this? x is 40, y is 84

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

I need more constraints and I am trying to find the feasible region

jimthompson5910 (jim_thompson5910):

when you solve 16f + 2790 = 3430, you get 16f + 2790 = 3430 16f = 3430-2790 16f = 640 f = 640/16 f = 40

jimthompson5910 (jim_thompson5910):

Then you use this to find c c = 124 - f c = 124 - 40 c = 84

jimthompson5910 (jim_thompson5910):

I don't see why you need to add constraints

OpenStudy (anonymous):

Well, my objective function was 38.50x +22.50y=3430 and I am supposed to write restraints and find the feasible regions to graph it.

jimthompson5910 (jim_thompson5910):

did you see how I got f = 40 and c = 84? those values mean that we have 40 full-size bags and 84 collapsible bags

OpenStudy (anonymous):

yes but how would I graph this to find the feasible region

jimthompson5910 (jim_thompson5910):

check: 40+84 = 124 38.50*40 + 22.50*84 = 3430 so each works out

jimthompson5910 (jim_thompson5910):

are you sure it doesn't say "up to 124 bags"? or anything like that?

OpenStudy (anonymous):

No I wrote the problem exactly as it says. I think I got it. I graphed a line for 124 and the 84 and 40 were on the line.

jimthompson5910 (jim_thompson5910):

honestly I'd just forget about the constraints and end the problem with saying "40 full-size bags and 84 collapsible bags" this is because constraints are usually for optimization problems

OpenStudy (anonymous):

I'd love to but I had to list the constraints and graph them and write a complete sentence about this problem. Love the common core.... I couldn't have done the problem without your help with the objective function, so thank you very much:)

OpenStudy (anonymous):

I'm new to Open Study, how to I give you a vote? other than just "best response?"

jimthompson5910 (jim_thompson5910):

hmm I see. I'm not familiar with the common core, but it sounds a bit frustrating. you click "best response" next to the person who you want to give a medal to

jimthompson5910 (jim_thompson5910):

i don't mind really, so either way is fine

OpenStudy (anonymous):

Ok, well, I did click, so I do really appreciate your help tonight:)

jimthompson5910 (jim_thompson5910):

you're welcome